WHAT IS THE SLOPE OF 3Y=X+3?
Found 2 solutions by ewatrrr, algebrahouse.com:
Answer by ewatrrr(24785)


Hi
Using the standard slope-intercept form for an equation of a line y = mx + b
where m is the slope and b the y-intercept.
3y = x+3
y = (1/3)x + 1 |Slope m = (1/3)

Answer by algebrahouse.com(1659)

WHAT IS THE SLOPE OF 3Y=X+3?

Slope-intercept form is y = mx + b
m is the slope
b is the y-intercept

3y = x + 3
y = (1/3)x + 1 {divided both sides by 3}

slope = 1/3
www.algebrahouse.com